OpenStudy (anonymous):

How do yo change how a fraction looks. Something about -1. Cant remember. Any ideas?

OpenStudy (anonymous):

You can change a fraction into a mixed fraction or improper fraction. A mixed fraction is something like this \[4\frac{2}{3}\] A improper fraction is something like this. \[\frac{14}{3}\]

OpenStudy (anonymous):

If you want to make an integer a fraction you just put 1 as the denominator... So if you want to make 5 a fraction. \[\frac{5}{1}\]

OpenStudy (anonymous):

This is not what i meant. I want to turn (4x-1) into 1-4x

OpenStudy (anonymous):

That's not a fraction.

OpenStudy (anonymous):

24x^4-8x^3+3/(4x-1)^2

OpenStudy (anonymous):

I wnat to change it into 24x^4-8x^3+3/(1-4x)^2

OpenStudy (anonymous):

why do you want to do that?

OpenStudy (anonymous):

Are you meant to simplify that fraction?

OpenStudy (anonymous):

Alternative form forgot how to do it

OpenStudy (anonymous):

If i divide the whole expression by -1/1 that works.

OpenStudy (anonymous):

but you can't change the inside when there's a power beside it...

OpenStudy (anonymous):

if it was (4x-1) then okay you can do it. But if it's (4x-1)^2, would only be changing one of the (4x-1). You would have this (1-4x)(4x-1)

OpenStudy (anonymous):

So whats the thinking behind changing it. Dont see how its been done.

OpenStudy (stamp):

Oh my goodness. This is the third time I have answered this question for you. \[(4x-1)^2=(4x-1)(4x-1)\]\[(4x-1)=-(1-4x)\]\[-(1-4x)*-(1-4x)=(-1)^2(1-4x)^2\]\[(4x-1)^2=(1-4x)^2\]